votecity20:

In a system of pulleys, the effort P required to raise a load W is given by P = aW + b, where a and b are constants. If W = 40 when P = 12 and W = 90 when P = 22, find the values of a and b

AZ:

if W = 40 when P = 12 and W = 90 when P = 22 what are the values of a and b? so the equation is: P = aW + b let's substitute those numbers in equation 1: W = 40 when P = 12 12 = 40a + b equation 2: W = 90 when P = 22 22 = 90a + b

AZ:

We have two equations 12 = 40a + b 22 = 90a + b Do you know how to solve a system of equations?

AZ:

You can either graph the two lines and find the point where they intersect. Or you can use either the elimination or substitution method to solve for the variables. Do you understand so far and know how to proceed now?

votecity20:

Yes Thank you